OpenStudy (anonymous):

x=7-y find the solution set if x 0 1 7 then y will be???

OpenStudy (anonymous):

x=7-y y=7-x Now substitute the values of x and get the value of y

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

not sure..must I try to cancel out x and y

OpenStudy (anonymous):

Nope, for example when x= 0 then y=7-0 = 7

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ok so it should look like this 7-0= 7 7-1=6 7-7=0

OpenStudy (anonymous):

Yep.....

OpenStudy (anonymous):

is that it

OpenStudy (anonymous):

yes thats it...... u got the solution set

OpenStudy (anonymous):

so how do i place the equation on a graph

OpenStudy (anonymous):

PLot the points (0,7) , (1,6) and (7,0) and goin with a straight line as it is a linear equation

OpenStudy (anonymous):

Oh..I get it..

OpenStudy (anonymous):

gee thanks a bunch

OpenStudy (anonymous):

Welcome